OpenStudy (anonymous):

give exact answer: 5(e)^x+1=2(3)^x

OpenStudy (anonymous):

5(e)x+1−2(3)x=0

OpenStudy (zzr0ck3r):

I don't think you are going to find a exact solution to this.

Directrix (directrix):

@yoitsren Would you post the problem exactly as it appears in your text? I think some misplaced or missing grouping symbols have messed up the problem you meant to post. You can click on the blue DRAW button below and write the problem if that helps.

OpenStudy (asnaseer):

if this is your question:\[5e^{x+1}=2\times3^x\]then first divide both sides by 5 to get:\[e^{x+1}=0.4\times3^x\]Then take logs of both sides to get:\[x+1=\ln(0.4\times3^x)=\ln(0.4)+\ln(3^x)=\ln(0.4)+x\ln(3)\]Hopefully you can solve for x from here...